OpenStudy (anonymous):

integral of (3x+1) dx? Upper limit =1. lower limit=0

OpenStudy (anonymous):

\[\int\limits_{0}^{1} (3x+1)dx \] please explain. Im getting the answer 4. But im pretty sure its not correct...

OpenStudy (anonymous):

(3x^2)/2 + x | (3(1)/2 + 1) -0 (because 3x0/2 +0 is obviously just 0) = 2 & 1/2

OpenStudy (anonymous):

woops. sorry, the equation is actually :: \[ \int\limits_{0}^{1} (3x+1)^{2} \]. The answer is 7 but im getting 4

OpenStudy (anonymous):

first, foil. (3x+1)^2= 9x^2+6x+1 dx = 3x^2+3x+x | (3+3+1) - 0 = 7

OpenStudy (anonymous):

I expanded (3x +1)^2 to \[(3x+1) (3x+1) \] Then factored it. Then integrated it ..

OpenStudy (anonymous):

okay, thank you!